An aircraft can fly 960 miles in 4 hours with the wind and travel the same distance against the wind in 6 hours. What is the wind speed?

Answer by stanbon(48551)

): An aircraft can fly 960 miles in 4 hours with the wind and travel the same distance against the wind in 6 hours. What is the wind speed?
-------------------
let "p" be the speed of the plane in calm air
Let "w" be the wind speed
-------------------
With the wind DATA:
distance = 960 miles ; time = 4 hrs ; rate = 960/4= 240 mph
-------------------
Against the wind DATA:
distance = 960 miles ; time = 6 hrs ; rate = 960/6 = 160 mph
-------------------
Eqations:
p + w = 240 mph
p - w = 160 mph
-------------------
2p = 400 mph
p = 200 mph (speed of the plane in calm air)
-------------------
200 + w = 240
w = 40 mph (wind speed)
========================
Cheers,
Stan H.